OpenStudy (anonymous):

Find the lateral area, total area, and volume for the cylinder with the given measurement. r = 2, h = 4 L.A. = , T.A. = , V = ,

OpenStudy (anonymous):

T.A. = 2(pi)r^2+2(pi)rh=8pi+16pi=24pi V=(pi)h*r^2=16pi i dont kno what L.A. is